for what values of x, y, z and r is the expression Mx^ ly^ T2^ Qr (i) Boyle’s law (ii) charles’s law (iii) pressure law (iv) general gas law
1
Expert's answer
2018-04-02T10:47:08-0400
(i)
x=1,y=2,z=-2,r=0.
(ii)
x=0,y=3,z=0,r=-1.
(iii)
x=1,y=-1,z=-2,r=-1.
(iv)
x=1,y=2,z=-2,r=-1.